# LLaRA: Supercharging Robot Learning Data for Vision-Language Policy

### v1

@@ Abstract
Large Language Models (LLMs) equipped with extensive world knowledge and strong reasoning skills can tackle diverse tasks across domains, often by posing them as conversation-style instruction-response pairs. In this paper, we propose LLaRA: Large Language and Robotics Assistant, a framework which formulates robot action policy as conversations, and provides improved responses when trained with auxiliary data that complements policy learning. LLMs with visual inputs, i.e., Vision Language Models (VLMs), have the capacity to process state information as visual-textual prompts and generate optimal policy decisions in text. To train such action policy VLMs, we first introduce an automated pipeline to generate diverse high-quality robotics instruction data from existing behavior cloning data. A VLM finetuned with the resulting collection of datasets based on a conversation-style formulation tailored for robotics tasks, can generate meaningful robot action policy decisions. Our experiments across multiple simulated and real-world environments demonstrate the state-of-the-art performance of the proposed LLaRA framework. The code, datasets, and pretrained models are available at https://github.com/LostXine/LLaRA.
Keywords: VLM, Robot Action Policy, Instruction Tuning Data
Figure 1: LLaRA Overview. We convert expert trajectories into instruction-tuning data, while also generating auxiliary data that complements policy in a self-supervised fashion. Based on such, we finetune a pre-trained Vision Language Model (VLM) as a robot policy. At inference, we query the VLM to generate actions in natural language, just as in a conversation. A real-world inference example is shown in the figure.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 [1], Llama [2, 3], and Gemini [4] exhibit unprecedented abilities across diverse language tasks. Such LLMs are often trained together with visual data by taking advantage of pretrained vision encoders, forming powerful Vision Language Models (VLMs). Effective and efficient training of such VLMs significantly depends on the style and format of the vision-language data used for training, leading to the exploration of various instruction tuning [5, 6] strategies. In fact, visual instruction tuning, or formatting image-text data in conversation style for VLM training has been studied extensively in [7, 8]. The resulting VLMs exhibit strong vision-language skills, but not without limitations such as spatial awareness [9, 10] or niche-domain understanding [11, 12]. The latter has led to domain-specific visual instruction tuning [11, 12, 13].
@@
Several robot learning approaches also leverage pretrained LLMs / VLMs [14, 15, 16, 17, 18, 19] with more-recent work focusing on low-level robot actions [20, 21, 22]. The strong performance of such approaches can be attributed to the extensive world knowledge (e.g., understanding of physics, human common-sense) [23, 24] and powerful reasoning abilities [25, 26] of underlying language models. However, studies on curating conversation-style data and using such for instruction tuning in robotics have been rather limited [27].
@@
Motivated by the promising attributes of VLMs, we explore a formulation of VLM-based robot action policy in this paper. Our goal is to adapt a VLM as a robot action policy that can handle diverse visuomotor control challenges, a process we call Visuomotor Instruction Tuning. More specifically, we transform a typical behavior cloning (BC) dataset into an instruction dataset, Instruct-BC (inBC), which is then used to finetune a VLM. Given a state described in visual-textual modalities, our VLM is trained to generate suitable actions as text. Such a formulation based on conversation-style instruction-response data enables us to instruction-tune and convert a VLM into a robot action policy effortlessly. However, we also note that the effectiveness of an instruction-tuned VLM depends heavily on the quality of data formulation [7, 10, 12], which is non-trivial to automate (or, scale) for new domains  [11, 12, 13, 9, 10]. Therefore, to strengthen (or, supercharge) such data, we construct auxiliary datasets that complement policy learning from the same BC dataset, in a self-supervised fashion without any external data.
@@
Our overall framework termed LLaRA (Large Language and Robotics Assistant) generates visuomotor instruction data that can efficiently and effectively finetune a VLM into a robot action policy. It is motivated by LLaVA [7] which is designed primarily for vision tasks. Similarly, LLaRA offers the entire framework of data generation, model formulation, and training pipeline for VLMs, now specialized for robot learning. We summarize our key contributions as follows:
@@
    A scalable pipeline for generating diverse high-quality robot instruction tuning data from existing behavior cloning data.
@@ 2 Related Work
Spatial Reasoning in VLMs: Several recent work investigate how VLMs can be modified for spatial awareness within images [28, 29, 30, 31, 9, 10, 32, 33, 34], which is a critical ability of a visuomotor policy. One line of work explores prompting using textual or specialized tokens to encode locations, followed by instruction tuning on localization-specific datasets  [31, 9, 10, 32]. In fact, several works [9, 10, 32] combine LLMs / VLMs with templating operations to generate datasets in formats adhering to human conversation, using existing annotated datasets that only contain per-object bounding box annotations. In contrast, our framework extends beyond localization to directly predict robot actions. An alternate line of work explores visual prompting but without any focus on robot learning (e.g., red circles [34] or free-form visual scribbles [33] overlaid on images images to focus a VLM on specific regions). An extension of such ideas to robotics is explored in PIVOT [35], where images annotated with object locations and trajectories are given to a VLM for generating visually-described robot actions. These ideas are complementary to LLaRA, where additional visuomotor instruction tuning directly generates actions optimal for robotics tasks.
Agent and Policy Learning with LLMs / VLMs: Foundation LLMs / VLMs provide a source of common-sense data for robot policy training. Qian et al. [36] shows how to train an affordance grounding VLM on affordance data. Ingelhag et al. [37] presents a system that evaluates the potential of instruction following using already-learned skills based on LLM / VLM, and otherwise learns new skills from user demonstrations. Wu et al. [38] construct a read-and-reward framework that extracts relevant information from Atari manuals, evaluating object-agent interactions. Yoneda et al. [39] utilizes an LLM to maintain an estimate of the world state in text. More-recent work such as Robotics Transformer [20, 21, 22], Gato [40], GR-1 [41] and Octo [42] learn generalist robot policies with multi-modal sequence models, taking advantage of the flexibility of encoding and decoding images / actions via token representations similar to VLMs. VIMA [43] is a transformer-based agent that processes interleaved visual and text prompts. In contrast, our method predicts actions encoded only as texts, without introducing any new tokens or LLM / VLM architectural modifications. LLaRA uniquely reformulates the same data in VIMA [43] for instruction tuning.
Instruction Data Generation: Finetuning LLMs / VLMs with instruction data has shown a lot of potential for vision tasks [5, 6, 7, 8]. While effective, construction of such datasets with high quality and quantity is beyond trivial [7], especially for specialized visual domains [11, 44, 12, 45] such as robotics. Moreover, recent works highlight shortcomings in abilities such as spatial awareness of VLMs instruction tuned on generic data [9, 10] leading to exploration of localization-specific instruction data generation [31, 9, 10, 32, 8]. Yet these modified VLMs suffer in the robotics domains [46]. In contrast to these works, we build a robotics-specific instruction dataset used to train a powerful VLM, LLaRA, adept at several robotic domain tasks. Closest to our proposed LLaRA is Zhen et al. [27], which focused on 3D reasoning and inference capabilities for manipulation tasks, but unlike ours perform no direct action predictions in textual form.
Self-Supervised Learning in Robotics: Self-supervised learning (SSL) has long been explored in robotics [47, 48], with recent work showing a flavor of visual representation learning  [49, 50, 51]. In contrast, we introduce auxiliary SSL objectives (i.e., pretext tasks) based on instruction prompts, motivated by similar ideas in computer vision (e.g., utilizing spatial relations in [52], temporal relations in [53, 54] or future prediction in [55]). More specifically, we create a set of domain-specific auxiliary datasets (e.g., for spatial or temporal reasoning), that enable LLaRA to learn relations that are useful for the downstream robotics tasks, when used for finetuning.
Discussion on Concurrent Work: Within the past 2 weeks prior to the release of our work, there has been a number of concurrent work studying VLM-based action policy learning. They have similarities with LLaRA, as they promote instruction tuning, motivated by prior work such as LLaVA [7]. Our key difference is the automated generation of auxiliary data from existing robot trajectories, that enables robots to learn spatio-temporal relations in the scene. We further discuss them below:
@@
All such approaches employ a VLM as a robot action policy, processing both visual observations and textual instructions as inputs. OpenVLA [18] employs specialized tokens to represent quantized actions, a method akin to RT-2 [21]. Additionally, OpenVLA incorporates DINOv2 [56] as an additional visual encoder along with SigLIP [57]. This combination has proven beneficial, particularly when the visual encoder is finetuned for robotics applications. RoboPoint [58] introduces a novel point-based action space and a scalable data pipeline tailored for spatial affordance prediction. Actions are delineated via points on an RGB image, which are subsequently translated to 3D space using depth data. This approach eliminates the reliance on predefined action primitives [59, 60], external object detectors [61, 62], and iterative visual prompting [35]. Another concurrent work, LLARVA [19], differs by generating both 2D visual traces in image coordinates and corresponding textual actions as outputs, with the former functioning as an auxiliary task.
However, a significant gap shown in the aforementioned studies is the lack of comprehensive investigation into the methods for generating auxiliary datasets from existing robot data, as well as the implications of integrating such datasets. Our work LLaRA distinctively addresses this gap, marking a critical divergence from the concurrent studies. That is, in addition to directly supervising a VLM based on robot actions, LLaRA framework automatically generate multiple auxiliary datasets with complementary training objectives (from the existing robot trajectories). Such auxiliary datasets are designed to train a VLM to capture spatial/temporal relations between objects in the scene, enabling better scene representations. We experimentally confirm that such additional supervision significantly benefits robot action policy learning.
@@ 3 Preliminary: Visual Instruction Tuning
Visual Instruction Tuning is a framework to finetune a large Vision Language Model (VLM) using language as task instructions. The objective is to develop a general-purpose multimodal model capable of learning from diverse vision tasks described by language [63]. A typical example of Visual Instruction Tuning is LLaVA [7], which consists of three neural networks as its model: a pretrained LLM $\theta_{\text{LLM}}$, a pretrained visual encoder $\theta_{V}$ and an adapter layer $\theta_{\text{MLP}}$.
@@
Figure 2: Background: LLaVA Overview. A Large Language Model (LLM) is connected to the visual domain with suitable encoder and adaptor neural networks.
A two-phase training process is executed in Liu et al. [7] with a next-token prediction objective over the $\theta_{\text{LLM}}$ output in each phase. First, only $\theta_{\text{MLP}}$ is trained using image-caption pairs (similar to datasets for CLIP [64] training). Then, both $\theta_{\text{MLP}}$ and $\theta_{\text{LLM}}$ are finetuned using a specialized instruction tuning dataset termed LLaVA_Instruct_150k (see Fig. 2). LLaVA_Instruct_150k contains human-style instruction-response conversations, automatically created by a powerful LLM [1] from a generic object detection dataset (see Liu et al. [7] for details). In our paper, we start from a pretrained LLaVA model and only perform a single-stage finetune, updating $\theta_{\text{LLM}}$ and $\theta_{\text{MLP}}$.
@@ 4 Visuomotor Instruction Tuning
We leverage large Vision Language Models (VLMs) as a generalist to address diverse visuomotor control challenges, a process we call Visuomotor Instruction Tuning. Specifically, we transform a typical behavior cloning dataset into an instruction tuning dataset and subsequently finetune a pretrained VLM on this tailored dataset. The resulting LLaRA framework benefits from the broad, inherent knowledge embedded within the VLM, enabling better visuomotor task learning.
@@
In this section, we present our LLaRA framework to a) convert a set of robot manipulation expert trajectories into a visuomotor instruction tuning dataset, and b) turn a VLM finetuned on such a dataset into a competitive robot policy. In the next section, we show that such a VLM can be further improved for robot control in a self-supervised manner using auxiliary instruction tuning datasets, which are also extracted from the same existing expert demonstrations.
4.1 Problem Formulation
@@
We consider a behavioral cloning (BC) setting over a Markov Decision Process (MDP), described by the tuple $(\mathcal{S},\mathcal{A},P)$. Here, $s_{t}\in\mathcal{S}$ denotes the state at timestamp $t$, $a_{t}\in\mathcal{A}$ represents the action at $t$, and $P$ encapsulates the transition dynamics, expressed as $P(s_{t+1}|s_{t},a_{t})$. Our goal is to finetune a VLM as a robot policy $\pi$ that best recovers an unknown policy $\pi^{*}$ using a demonstration dataset containing multiple tasks $\mathcal{D}=\{(s^{i},a^{i})\}$ collected by $\pi^{*}$. The robotic policy $\pi$ is designed to be conditioned on the task description, and our empirical findings suggest adding historical actions is a beneficial condition. Consequently, the policy can be articulated as $\pi(a_{t}^{i}|s_{t}^{i},a_{h}^{i})$, where $h=1,2,...,t-1$, indicating that $\pi$ takes into account all preceding actions up to time $t-1$.

### v2

@@ Abstract
LLMs with visual inputs, i.e., Vision Language Models (VLMs), have the capacity to process state information as visual-textual prompts and respond with policy decisions in text. We propose LLaRA: Large Language and Robotics Assistant, a framework that formulates robot action policy as conversations and provides improved action outputs when trained with auxiliary data that complements policy learning. We first introduce an automated pipeline to generate conversation-style instruction tuning data from existing behavior cloning data. Then we enrich the dataset in a self-supervised fashion by formulating six auxiliary tasks. A VLM finetuned with the resulting collection of datasets can generate meaningful robot action policy decisions. Our experiments across multiple simulated and real-world environments demonstrate the state-of-the-art performance of the proposed LLaRA framework. The code, datasets, and pretrained models are available at https://github.com/LostXine/LLaRA.
Figure 1: A real-world demonstration of LLaRA solving an unseen task. In the training stage, expert trajectories are converted into instruction-tuning data with image coordinates. Additionally, auxiliary data is generated to enhance the policy in a self-supervised manner. Using this data, we finetune a pretrained Vision Language Model (VLM) as a robot policy. At inference, the VLM is queried to generate actions in natural language.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), Llama (Touvron et al., 2023; AI, 2024), and Gemini (Anil et al., 2023) exhibit unprecedented abilities across diverse language tasks. Such LLMs are often trained together with visual data by taking advantage of pretrained vision encoders, forming powerful Vision Language Models (VLMs). Effective and efficient training of such VLMs significantly depends on the style and format of the vision-language data used for training, leading to the exploration of various instruction tuning (Taori et al., 2023; Chiang et al., 2023; Liu et al., 2023a; 2024b) strategies. The resulting VLMs exhibit strong vision-language skills, but not without limitations such as spatial awareness (Chen et al., 2023; Ranasinghe et al., 2024b) or niche-domain understanding. The latter has led to domain-specific visual instruction tuning (Li et al., 2023; Kuckreja et al., 2024; Hong et al., 2023).
@@
Several robot learning approaches also leverage pretrained LLMs / VLMs (Zeng et al., 2022; Brohan et al., 2023c; Driess et al., 2023; Vemprala et al., 2024; Kim et al., 2024; Niu et al., 2024) with recent work focusing on low-level robot actions (Brohan et al., 2023b; a; Padalkar et al., 2023). The strong performance of such approaches can be attributed to the extensive world knowledge (e.g., understanding of physics, human common-sense) (Yu et al., 2023; Zhao et al., 2024) and powerful reasoning abilities (Creswell & Shanahan, 2022; Liu et al., 2023b) of underlying language models. However, studies on curating and using conversation-style data for instruction tuning in robotics have been rather limited (Zhen et al., 2024).
@@
Motivated by the promising attributes of VLMs, we explore a formulation of VLM-based robot action policy in this paper. Our goal is to adapt a VLM as a robot action policy that can handle diverse visuomotor control challenges, a process we call Visuomotor Instruction Tuning. More specifically, we transform a typical behavior cloning (BC) dataset into an instruction dataset, Instruct-BC (inBC), which is then used to finetune a VLM. Given a state described in visual-textual modalities, our VLM is trained to generate suitable actions as text. Such a formulation based on conversation-style instruction-response data enables us to instruction-tune and convert a VLM into a robot action policy effortlessly. However, we also note that the effectiveness of an instruction-tuned VLM depends heavily on the quality of data formulation (Liu et al., 2023a; Ranasinghe et al., 2024b; Kuckreja et al., 2024), which is non-trivial to automate (or, scale) for new domains  (Li et al., 2023; Kuckreja et al., 2024; Hong et al., 2023; Chen et al., 2023; Ranasinghe et al., 2024b). Therefore, to strengthen (or, supercharge) such data, we construct auxiliary datasets that complement policy learning from the same BC dataset, in a self-supervised fashion without any extra action label.
@@
Our overall framework termed LLaRA (Large Language and Robotics Assistant) generates visuomotor instruction data that can efficiently and effectively finetune a VLM into a robot action policy. It is motivated by LLaVA (Liu et al., 2023a) which is designed primarily for vision tasks. Similarly, LLaRA offers the entire framework of data generation, model formulation, and training pipeline for VLMs, now specialized for robot learning. Our key contributions are:
@@
    Aligning robot actions to image coordinates for VLMs, making the transfer of a pretrained VLM to the robotics domain easier, particularly when the training data is limited.
@@ 2 Related work
Instruction data generation. Finetuning LLMs / VLMs with instruction data has shown a lot of potential for vision tasks (Taori et al., 2023; Chiang et al., 2023; Liu et al., 2023a; 2024b). While effective construction of such datasets with high quality and quantity is beyond trivial (Liu et al., 2023a), especially for specialized visual domains (Li et al., 2023; Bazi et al., 2024; Kuckreja et al., 2024; Thawkar et al., 2023) such as robotics. Yet these modified VLMs suffer in the robotics domains (Ranasinghe et al., 2024a).
@@
Spatial reasoning in VLMs. Several recent works investigate how VLMs can be modified for spatial awareness within images (Zhang et al., 2023; Zhao et al., 2023; Zang et al., 2023; Peng et al., 2023; Chen et al., 2023; Ranasinghe et al., 2024b; You et al., 2023; Cai et al., 2024; Shtedritski et al., 2023), which is a critical ability of a visuomotor policy. One line of work explores prompting using textual or specialized tokens to encode locations, followed by instruction tuning on localization-specific datasets  (Peng et al., 2023; Chen et al., 2023; Ranasinghe et al., 2024b; You et al., 2023). In contrast, our framework extends beyond localization to directly predict robot actions. An alternate line of work explores visual prompting but without any focus on robot learning  (Shtedritski et al., 2023; Cai et al., 2024). An extension of such ideas to robotics is explored in PIVOT (Nasiriany et al., 2024). These ideas are complementary to LLaRA, where additional visuomotor instruction tuning directly generates actions optimal for robotics tasks.
Robot learning via sequence models, LLMs, and VLMs. The use of sequence models in robot learning has a longstanding tradition (Zheng et al., 2022; Shang et al., 2022a; b). Recent foundation LLMs / VLMs provide a source of common-sense data for robot policy training (Qian et al., 2024; Ingelhag et al., 2024; Wu et al., 2023b; Yoneda et al., 2023). More recent work such as Robotics Transformer (Brohan et al., 2023b; a; Padalkar et al., 2023), Gato (Reed et al., 2022), GR-1 (Wu et al., 2023a) and Octo (Octo Model Team et al., 2024) learn generalist robot policies with multi-modal sequence models, taking advantage of the flexibility of encoding and decoding images/actions via token representations similar to VLMs.
PALM-E (Driess et al., 2023) utilizes a pretrained VLM for embodied reasoning and planning. However, it assumes access to additional policies capable of performing low-level skills from a limited vocabulary, with the VLM generating natural language that conditions these low-level commands. Our method requires neither a language-conditioned low-level controller, which can be expensive to learn for each environment, nor a lot of training data to achieve good performance. In addition, because we use 2D image coordinates to present position, one can easily transfer our model to another embodiment/domain with limited training samples. SpatialVLM (Chen et al., 2024) focuses on enhancing the spatial reasoning ability of the VLM, resulting in a VLM that serves as a spatial relationship predictor rather than a robot policy. To embed SpatialVLM into a policy, chain-of-thought spatial reasoning is required. Our method explicitly transforms a pretrained VLM into a robot policy that directly generates robot actions, resulting in a simpler framework and no need for an additional policy. In addition, SpatialVLM studies only the spatial relationships between objects. Our method also learns temporal environment dynamics through action prediction and future prediction datasets, making our VLM better prepared as a robot policy.
@@
There are concurrent works (Kim et al., 2024; Yuan et al., 2024; Niu et al., 2024) which also utilize a LLaVA style VLM as a robot policy. All such approaches employ a VLM as a robot action policy, processing both visual observations and textual instructions as inputs. OpenVLA (Kim et al., 2024) employs specialized tokens to represent quantized actions, a method akin to RT-2 (Brohan et al., 2023a). OpenVLA also incorporates DINOv2 (Oquab et al., 2023) as an additional visual encoder along with SigLIP (Zhai et al., 2023), compared to a single CLIP (Radford et al., 2021) vision encoder used in LLaVA. RoboPoint (Yuan et al., 2024) introduces a novel point-based action space and a scalable data pipeline tailored for spatial affordance prediction. Actions are delineated via points on an RGB image, which are subsequently translated to 3D space using depth data. This approach eliminates the reliance on predefined action primitives (Liang et al., 2023; Singh et al., 2023), external object detectors (Huang et al., 2023b; Liu et al., 2024a), and iterative visual prompting (Nasiriany et al., 2024). Another concurrent work, LLARVA (Niu et al., 2024), differs by generating both 2D visual traces in image coordinates and corresponding textual actions as outputs, with the former functioning as an auxiliary task. In contrast, our method describes robot action using image coordinates in natural language, which creates a more intuitive link between vision and action without modifying the existing VLM model architecture. Moreover, all the aforementioned studies lack of comprehensive investigation into the methods for generating auxiliary datasets from existing robot data, as well as the implications of integrating such datasets. Our work LLaRA distinctively addresses this gap, marking a critical divergence from the concurrent studies. That is, in addition to directly supervising a VLM based on robot actions, the LLaRA framework automatically generates multiple auxiliary datasets with complementary training objectives from the existing robot trajectories. Such auxiliary datasets are designed to train a VLM to capture spatial/temporal relations between objects in the scene, enabling better scene understanding. We experimentally confirm that such additional supervision significantly benefits robot action policy learning.
Self-supervised learning in robotics. Self-supervised learning (SSL) has long been explored in robotics (Pinto & Gupta, 2016; Sermanet et al., 2018; Yarats et al., 2021; Laskin et al., 2020; Schwarzer et al., 2021), with recent work showing a flavor of visual representation learning  (Li et al., 2022; 2024b; Wang et al., 2023). We introduce auxiliary SSL objectives (i.e., pretext tasks) based on instruction prompts, motivated by similar ideas in computer vision (Doersch et al., 2015; Yun et al., 2022; Ranasinghe et al., 2022; Walker et al., 2021). More specifically, we create a set of domain-specific auxiliary datasets (e.g., action or future prediction), that enable LLaRA to understand information that is useful for the downstream robotics tasks, when used for finetuning.
@@ 3 Preliminary: visual instruction tuning
Figure 2: LLaVA overview. A Large Language Model (LLM) is connected to the visual domain with suitable encoder and adaptor neural networks.
Visual Instruction Tuning is a framework to finetune a large Vision Language Model (VLM) using task-specific instructions. The objective is to develop a general-purpose multimodal model capable of solving diverse vision tasks described by language (Huang et al., 2023a). A typical example of Visual Instruction Tuning is LLaVA (Liu et al., 2023a), which consists of three neural networks as its model: a large language model $\theta_{\text{LLM}}$, a visual encoder $\theta_{V}$ and an adapter layer $\theta_{\text{MLP}}$ (see Fig. 2).
@@
A two-phase training process is executed in Liu et al. (2023a) with a next-token prediction objective over the $\theta_{\text{LLM}}$ output in each phase. First, only $\theta_{\text{MLP}}$ is trained using image-caption pairs (similar to datasets for CLIP (Radford et al., 2021) training). Then, both $\theta_{\text{MLP}}$ and $\theta_{\text{LLM}}$ are finetuned using a specialized instruction tuning dataset termed LLaVA_Instruct_150k. LLaVA_Instruct_150k contains human-style instruction-response conversations, automatically created by a powerful LLM (OpenAI, 2023) from a generic object detection dataset; see (Liu et al., 2023a) for details. In our paper, we start from a pretrained LLaVA model and only perform a single-stage finetune, updating $\theta_{\text{LLM}}$ and $\theta_{\text{MLP}}$.
@@ 4 Visuomotor instruction tuning
We leverage large Vision Language Models (VLMs) as a generalist to address diverse visuomotor control challenges. Specifically, we transform a typical behavior cloning dataset into an instruction tuning dataset and subsequently finetune a pretrained VLM on this tailored dataset, a process we call Visuomotor Instruction Tuning. The resulting LLaRA framework benefits from the broad, inherent knowledge embedded within the VLM, enabling better visuomotor task learning.
@@
In this section, we present our LLaRA framework to a) convert a set of robot manipulation expert trajectories into a visuomotor instruction tuning dataset, and b) finetune a VLM on such a dataset as a robot policy. In the next section, we show that c) such VLMs can be further improved for robot control using auxiliary instruction tuning data created in a self-supervised manner.
@@
Problem formulation. We consider a behavioral cloning (BC) setting over a Markov Decision Process (MDP), described by the tuple $(\mathcal{S},\mathcal{A},P)$. Here, $s_{t}\in\mathcal{S}$ denotes the state at timestamp $t$, $a_{t}\in\mathcal{A}$ represents the action at $t$, and $P$ encapsulates the transition dynamics, expressed as $P(s_{t+1}|s_{t},a_{t})$. Our goal is to finetune a VLM as a robot policy $\pi$ that best recovers an unknown policy $\pi^{*}$ using a demonstration dataset containing multiple tasks $\mathcal{D}=\{(s^{i},a^{i})\}$ collected by $\pi^{*}$. The policy $\pi$ is designed to be conditioned on the task description, and our empirical findings suggest adding historical actions is a beneficial condition. Consequently, the policy can be articulated as $\pi(a_{t}^{i}|s_{t}^{i},a_{h}^{i})$, where $h=1,2,...,t-1$, indicating that $\pi$ takes into account all preceding actions.
Instruction tuning data from trajectories. Next, we turn a set of expert trajectories $\mathcal{D}$ into a visuomotor instruction tuning dataset. We focus on the stationary robot manipulation scenario featuring a robotic manipulator fixed above a flat table. The visual observation for this setup is captured by a fixed third-person view camera positioned to observe the table surface. For each state transition, we convert the state action pair $(s_{t},a_{t})$ into a single-round conversation. The current visual observation and textual task description, forming $s_{t}$, can be directly assimilated by a VLM as the user instruction. However, the numerical action $a_{t}$ needs conversion into textual format to be generated by a VLM. In contrast to the approach employed by RT-2 (Brohan et al., 2023a), which utilizes special tokens to directly encode numerical action values, we adopt 2D Image Coordinates—normalized to $[0,1]$—to represent the positions in actions. Additionally, any rotational component of the action, denoted as $a_{r}$ (such as the angle of a joint), is expressed textually as ‘‘rotate <r>$a_{r}$</r> degrees’’. To further improve the performance, we incorporate the history of executed actions into the query and structure the response to predict multiple future action steps until the episode concludes. More details are covered in Appendix A.1.
Our approach creates an intuitive connection between visual input and robotic actions, enabling effortless transfer across various robotic embodiments. By implementing these strategies, we effectively convert a trajectory into a format comprising both image and text, which is readily processed by the VLM. We name the converted dataset Instruct-BC (inBC) and present examples in Fig. 3, Tab. 5 and Tab. 6.
Aggregate multiple images for a single-image VLM. The VLM policy we use here closely follows the design of LLaVA-1.5 (Liu et al., 2024b). The model consumes a single image and language instruction and generates language output. In scenarios where a single observation $s_{t}$ comprises multiple images, systems such as LLaVA suffer from low performance because it was never trained on multiple images (see Tab. 11). Instead, we convert each additional image into a language description with the help of object detection on the additional images (e.g., images in the task description, rather than the main visual observation). The main image is still directly used as the visual input by the VLM without any changes. Compared to inBC, we name this method that takes additional object detection results as Description-Instruct-BC ($\mathcal{D}$-inBC), and our model trained on this data is referred to with the same name as earlier. Examples are also available in Fig. 3, Tab. 5 and Tab. 6.
Inference pipeline. During inference, the prompt for the VLM is prepared using the same template as either inBC or $\mathcal{D}$-inBC. As illustrated in Fig. 3, for a model trained on inBC, each conversation turn encapsulates the current visual observation, the task description, and the previous actions described in the text. For a model trained on $\mathcal{D}$-inBC, the process starts with object detection to convert any extra images from the task description into text. The task description, now including the detection outcomes, is then combined with the current visual observation and the action history to get the complete query. Then LLaRA generates the text output that contains numerical values (e.g., 2D image coordinates) for a robot action. The extracted rotation angle can be directly mapped to the rotation of the end effector. The 2D image coordinates will be further converted to robot action space via predefined mapping, which can be estimated using visual calibration in both simulated and real-world environments. After that, the robot executes the action, and a new observation can be captured from the environment. This initiates another round of interaction, setting the stage for the subsequent action.